OpenStudy (anonymous):

Use the definitions sin0 y/r  , cos0 x/r,  and tan0 y/x  to prove each identity. (a) cos0/sin0 = 1/tan0

OpenStudy (kropot72):

 \[\frac{\cos \theta}{\sin \theta}=\frac{x}{r}\times \frac{r}{y}=\frac{x}{y}\] It is given that\[\tan \theta=\frac{y}{x}\] Can you finish the proof?